OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = –x + 7 A. slope: –7, y-intercept: 1 B. slope: –1, y-intercept: 7 C. slope: 1, y-intercept: –7 D. slope: 7, y-intercept: –1

OpenStudy (anonymous):

B

OpenStudy (anonymous):

slope formula, y = mx + b m is the slope, and b is the y-intercept. your equation is y = -x + 7 so m = -1 and b = 7 answer is B

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

no problem